# FLEX: A Large-Scale Multi-Modal Multi-Action Dataset for Fitness Action Quality Assessment

### v1

@@ 4.3 Novel Task–Video2EMG
sEMG provides feedback on muscle functional state and correlates with action quality. We envision a future where EMG signal could be estimated directly videos. Such technology would enable provide precise—muscle-level—feedback and intervention. To that end, we introduce a new task of estimating sEMG signals from videos. Our FLEX dataset with paired recordings of videos and EMG signals can enable this new research direction.
Video2EMG Baseline Model. Inspired by the emg2posesalter2024emg2pose framework, we designed an end-to-end visual regression model combining a ResNet backbone with an LSTM.
Implementation details. Provided in the Appendix. Code for this model has been released.
Metrics. To evaluate the Video2EMG model’s regression performance, we report mean absolute error (MAE) and root mean square error (RMSE), capturing average error and its variability, respectively.
Figure 8: Performance of Video2EMG models on FLEX dataset.
@@
Results. After 150 training epochs, the Video2EMG model exhibited strong regression performance (see in Figure 8). Following inverse normalization, it achieved an MAE of 0.1655 and an RMSE of 0.2134, demonstrating accurate sEMG value prediction across most samples while keeping error variability within a reasonable range. These results serve as baseline for future work.
5 Limitations and Future Work
Limitations. Despite our efforts to broaden recruitment, several contextual factors limited the inclusion of female participants in the FLEX dataset. The institute’s remote location, local gender imbalance, and sociocultural norms posed significant challenges. As a result, the dataset comprises 38 male participants. We acknowledge this limitation and are committed to improving gender diversity in future studies through expanded recruitment strategies and partnerships.
Future Work. FLEX dataset provides the computer vision community with novel, high-quality multimodal resources and establishes an extensible paradigm for future fitness datasets. Our preliminary VLM and Video2EMG experiments further demonstrate the promise of cross-modal research and integrating AQA models with vision–language models to realize a unified “action agent.”
@@ 6 Conclusion
In this paper, we propose FLEX, a large-scale, multimodal, multi-action, and comprehensive fitness action quality dataset. FLEX leverages a high-precision markerless motion capture system to record 20 weight-loaded fitness actions performed by 38 subjects across three ability levels, capturing RGB video from five viewpoints, 3D pose data, surface EMG signals, and physiological parameters. Moreover, FLEX is the first AQA dataset to integrate sEMG and a knowledge graph: by defining penalty-function-based annotation rules, we establish correspondences among weight-loaded actions, action knots, error types, and feedback. We conduct baseline experiments and novel applications on FLEX to demonstrate the effectiveness and potential of its multimodal data. We anticipate that FLEX will serve as a high-quality, extensible resource for researchers in the field.
Acknowledgements
This work is supported by the Suzhou Basic Scientific Research Project, Grant number SSD2024013. Fast Move Technology is especially acknowledged for providing sEMG devices, which greatly contributed to this research.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] Please refer to section 1.
    2. (b)
      Did you describe the limitations of your work? [Yes] Please refer to section 5.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] Please refer to the Appendix.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Please refer to the Appendix.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Please refer to the Appendix.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] We did not report error bars due to resource constraints, although we followed the standard evaluation setup commonly used in AQA studies.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Please refer to the Appendix.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] Please refer to section 4.
    2. (b)
      Did you mention the license of the assets? [Yes] Please refer to the Appendix.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] Please refer to the Appendix. And we got the permission from chair that we put the full version of the dataset on self self-hosted cloud drive.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] Please refer to section 3.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] Please refer to section 3.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] Please refer to the Appendix.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [Yes] Please refer to section 3
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] Please refer to the Appendix.
7 FLEX Dataset
7.1 Action/Exercise Selection
@@
We began by analyzing action types by conducting a comprehensive survey of publicly available fitness AQA datasets[52]. Our survey revealed that the Waseda-Squat[8] and Egoexo-Fitness[7] comprise a total of 12 self-loaded fitness actions, whereas the Fitness-AQA[6] includes 3 weight-loaded fitness actions. Self-loaded fitness actions do not rely on equipment, making them simpler and less prone to injury—thus more suitable for limited spaces or beginners. Conversely, weight-loaded fitness actions typically involve greater difficulty, with clearly defined target muscle groups and structured progression paths (e.g., systematically increasing dumbbell or barbell weights). Given the relatively narrow range of weight-loaded fitness actions in existing datasets, the FLEX dataset prioritizes these actions. However, due to time and resource constraints, collecting every possible weight-loaded fitness action was not feasible. Therefore, based on three dimensions—action prevalence, target muscle groups, and required equipment - we ultimately selected 20 common weight-loaded actions (including 10 barbell-based and 10 dumbbell-based actions). The chosen fitness actions not only closely mirror practical training regimens but also comprehensively cover both upper and lower body muscle groups, offering a more complete dataset compared to Egoexo-Fitness[7] (see in Figure 9 and Table 2). Additionally, researchers at Beijing Sport University validated the placement locations of the sEMG sensors on the target muscles for each action.
@@
Egoexo-Fitness[7] 1. Kneeling Push-ups 2. Push-ups 1. Pectoralis major 2. Anterior deltoid
@@ 13. Hip abductors
7.2 Subject Recruitment
Humans, as the core component in action performance, directly influence action quality. To collect more comprehensive data, the FLEX dataset required subjects across various capability levels compared with datasets that only contained professional-level subjects. So, we extensively recruited subjects within our institution and local commercial gyms, and ultimately selected 38 male subjects, comprising 10 professional coaches, 8 amateurs, and 20 novices. We entered into a data-collection agreement with each subject, securing their authorization to record video and physiological data and to make these publicly available for academic research; the agreement also explicitly delineates the applicable financial compensation terms. As noted above, to ensure data quality, only the professional group was required to complete all 20 action tests in one continuous session; the remaining 28 subjects were divided by target muscle groups and completed their tests incrementally over several days. Given the differences in testing intensity, we implemented a differentiated compensation scheme: professional subjects received ¥350 for completing all tests within approximately two hours, while the other subjects completed their sessions across three days, about 30 minutes per day, and were compensated ¥160 per day. (The minimum pay in Suzhou is ¥24/h.)
7.3 Collection Instruction
As noted above, prior to data collection, the experimenter outfitted each subject with a custom vest and had surface EMG sensors affixed to the target muscle regions. Subjects then entered the capture area individually, and upon the experimenter’s start signal, each subject performed the action ten times based on their own experience. Notably, the experimenter informed subjects of the specific action to be recorded before each trial, and the equipment load was set to 80% of the subject’s maximum to ensure that the action was executed by their own capability.
7.4 Annotation Rules
@@
Figure 10: The overview of the FLEX knowledge graph. (a) Visualization of frequently used annotation words. (b) FLEX-KG: the structure of the knowledge graph. (c1) Mapping between actions and action knots. (c2) Mapping between action knots and error types.
7.5 Annotator Recruitment
Due to the vast volume of the FLEX dataset and stringent annotation quality requirements, we recruited 16 professional practitioners from the fitness domain, including fitness trainers and graduate students in sports science, to participate in the annotation work. To ensure consistency and scientific rigor throughout the annotation process, we engaged in in-depth collaboration with all annotators before project commencement, clarifying the annotation requirements, personnel qualifications, and the underlying logic of the guidelines. To validate the feasibility and authority of these rules, we initially conducted a small-scale pilot annotation, during which we reviewed the annotators’ professional credentials and, based on the pilot results, made necessary revisions to the guidelines and refined the selection of annotators. The payment of a single annotator is higher than the minimum pay in Suzhou.
7.6 Score Calculation
During the annotation-rule formulation phase, we assigned different penalty weights to error types according to their impact on action quality. We then summed the weights of all error types present in a given action to obtain that action’s total weight. For each sample, based on the error types identified by the annotators, we calculated the ratio of the sample’s cumulative error weight to the action’s total weight, and subsequently computed the action score using the following formula (see in Figure 11):
@@ 7.6 Score Calculation
7.7 Data Statistics

### v3

@@ 4.3 Novel Task–Video2EMG
Figure 7: Qualitative result of Video2EMG. Notice the faithful prediction of EMG.
Surface-EMG (sEMG) signals provides crucial feedback on muscle functional state and correlates with action quality. However, acquiring sEMG data is costly. To mitigate that, we envision a future where EMG signal could be estimated directly from videos. Such technology would enable provide precise—muscle-level—feedback and intervention in cost-effective manner. To that end, we introduce a new task of estimating sEMG signals from videos. Our FLEX dataset with synchronized recordings of videos and EMG signals can enable this new research direction. Video2EMG Baseline Model: Inspired by the emg2pose Salter et al. (2024) framework, we designed end-to-end visual regression models that integrate a visual encoder (ResNet or ViT) with an sEMG sequence predictor (LSTM or stacked SVR). Implementation details provided in Appendix C. Code will be released.
Figure 8: Performance of Video2EMG models.
@@
ResNet-SVR 0.1466 0.1878
ViT-LSTM 0.1648 0.2069
ViT-SVR 0.1465 0.1882
Metrics. To evaluate Video2EMG regression, we report mean absolute error (MAE) and root mean square error (RMSE), reflecting average error and variability.
Results. After 150 training epochs, Video2EMG models exhibited strong regression performance (see in Figure 8). Following inverse normalization, they achieved the best MAE of 0.1465 and RMSE of 0.1882, demonstrating accurate sEMG value prediction across most samples while keeping error variability within a reasonable range. These results serve as baseline for future work.
@@ 5 Conclusion
We presented FLEX, a large-scale, multimodal dataset for fitness action quality assessment that pairs five-view RGB video, 3D pose, surface EMG, and physiological signals across diverse weight-loaded exercises. FLEX is the first AQA resource to integrate sEMG and a Fitness Knowledge Graph, providing structured annotations and penalty-based scoring that capture relationships among actions, key steps, errors, and corrective feedback. Baseline experiments—including multimodal AQA, VideoQA, and Video→EMG prediction—demonstrate the potential of FLEX for biomechanically grounded representation learning and structured reasoning. We hope FLEX will catalyze future research in multimodal fitness understanding, cross-modal prediction, and interpretable AQA research.
Ethics statement
Our work adheres to the conference’s Code of Ethics, emphasizing responsibility, fairness, transparency, and the minimization of harm in research and its applications.
Subjects Protection, Safety, and Privacy. To ensure the safety of participants during the collection of risk-prone exercise data, all subjects performed movements at approximately 80% of their maximum weight-loaded capacity. Safety supervisors were present on-site throughout the sessions, with emergency medical kits available. Additionally, all participants were covered by sports injury insurance. The research protocol was reviewed and approved by the Institutional Review Board (IRB), and written informed consent was obtained from all participants prior to data collection. To further protect participant privacy, all facial regions in the figures were blurred. Access to the dataset requires signing a License Agreement, which restricts usage to non-commercial academic research only, ensuring rigorous protection of personal data.
Fair and Responsible Compensation. Given the differences in testing intensity, we implemented a differentiated compensation scheme: expert subjects received compensation equivalent to roughly 15 times the local minimum hourly wage for completing all tests within approximately two hours, while other subjects, who completed their sessions across three days ($\sim$30 minutes per day), were compensated at a rate equivalent to about 7 times the local minimum hourly wage per day. Annotators were also compensated at a rate higher than the local minimum hourly wage, reflecting our commitment to fairness and ethical treatment of research contributors.
Fairness and Inclusivity. We issued an open call for participation in our data collection, welcoming individuals from diverse genders, ethnic groups, regions, and socioeconomic backgrounds. Participation from women was not observed in this phase of data collection. Contributing factors include the institute’s remote location (distant from urban centers), local demographic imbalances, and sociocultural dynamics. All recruitment followed established ethical principles in AI and data science, including informed consent, voluntary participation, non-coercive recruitment, and transparent communication of dataset limitations. To date, existing literature on AQA has not reported gender-specific performance differences in models. To continue strengthening inclusivity, we have initiated collaborations with partner institutions. Additional contributions from female subjects are already underway and are expected to be integrated into the dataset by the end of this year, further enriching its diversity and representativeness.
Awareness of Potential Negative Societal Impacts. The capability of measuring the quality of actions and movements may be used in ways that users/humans may not approve/agree. For example, action quality assessment may be misused by agencies such as health insurance providers to forecast future injuries and diseases reliably. Such health insurance providers may not then unjustly cover future injury-prone clientele. Awareness is the first step towards the solution. We believe our work helps create this awareness. Moreover, detecting bias of this nature in the decision-making process can help provide more equitable service and consideration to society.
Reproducibility statement
Due to the double-blind review policy of ICLR, we do not release the dataset or source code at this stage. Nevertheless, we have taken careful measures to ensure the reproducibility of our work. The main paper and Appendix provide detailed descriptions of the dataset construction process, including data collection protocols and preprocessing steps (see in the Sec. 3 and Appendix B). Comprehensive experimental details, including training configurations, hyperparameters, and evaluation protocols, are documented in the Experiments section (Sec. 4) of the main text and in the Appendix C. Together, these materials provide sufficient information for independent researchers to reproduce and verify our results. We plan to release the dataset and code publicly after the review process.
APPENDIX
Appendix A Related Work
VideoQA Dataset
In the VideoQA field, dataset development has gradually evolved from short-video scene understanding to multimodal spatiotemporal reasoning. Early datasets (e.g., MSVD-QA, MSRVTT-QA Xu et al. (2017), TGIF-QA Jang et al. (2017)) were based primarily on open-domain short clips, with questions centered on high-level semantics such as “What is someone doing?”, “What happened at a given moment?”, or “What is likely to happen next?”. Later datasets, including ActivityNet-QA Yu et al. (2019) and TVQA/TVQA+ Lei et al. (2019), introduced longer video segments, richer contextual information, and spatiotemporal localization annotations, advancing research in cross-modal alignment and temporal reasoning.
Despite this progress, human-action coverage in existing datasets remains coarse: most focus on broad categories (e.g., “running,” “jumping,” “lifting”) without fine-grained annotations of posture, deviations, or quality. Moreover, few—if any—datasets incorporate question–answering mechanisms with feedback or corrective functionality, such as “How should this be improved?” or “What adjustments should be made?”. In the AQA domain, datasets Parmar & Tran Morris (2017); Xu et al. (2022); Zhang et al. (2023) typically provide only scores or phase-level labels rather than QA-style supervision. Although Wu et al. Wu et al. (2025) attempted to augment existing AQA datasets with textual annotations, the lack of expert validation limited their reliability.
Compared with conventional VideoQA benchmarks, FLEX-VideoQA is tasked with a markedly different orientation: it not only requires models to recognize action categories and scene semantics but also emphasizes diagnosing posture and execution deviations, while further generating targeted improvement suggestions. This shift—from “What happened?” to “How should it be improved?”—underscores FLEX-VideoQA’s distinctiveness and innovative value within the broader VideoQA research landscape.
Muscle activity estimation work.
Some work like Peng et al. (2024) pursue predicting active muscle from video based on actions. However, there are essential differences in our task objectives and definition, and the fine-grained nature of our work as compared to active muscle detection. For example, these work tend to learn to correlate active muscle groups with coarse-grained actions, not necessarily ground it into actual muscle usage. In comparison, our work makes use of true muscle activity for the objective measuring action quality in action quality assessment task; or true estimation of muscle activity in the form of sEMG signals from videos, resulting in learning biomechanically-oriented representations.
Appendix B FLEX Dataset
B.1 Setup
B.1.1 Action/Exercise Selection
@@
We began by analyzing action types through a comprehensive survey of publicly available fitness AQA datasets, as summarized in a recent survey paper Yin et al. (2025). The survey revealed that the Waseda-Squat Ogata et al. (2019) and Egoexo-Fitness Li et al. (2024) comprise a total of 12 self-loaded fitness actions, whereas the Fitness-AQA Parmar et al. (2022) includes 3 weight-loaded fitness actions. Self-loaded fitness actions do not rely on equipment, making them simpler and less prone to injury—thus more suitable for limited spaces or beginners. Conversely, weight-loaded fitness actions typically involve greater difficulty, with clearly defined target muscle groups and structured progression paths (e.g., systematically increasing dumbbell or barbell weights). Given the relatively narrow range of weight-loaded fitness actions in existing datasets, the FLEX dataset prioritizes these actions. However, due to time and resource constraints, collecting every possible weight-loaded fitness action was not feasible. Therefore, based on three dimensions—action prevalence, target muscle groups, and required equipment — we ultimately selected 20 common weight-loaded actions (including 10 barbell-based and 10 dumbbell-based actions). The chosen fitness actions not only closely mirror practical training regimens but also comprehensively cover both upper and lower body muscle groups, offering a more complete dataset compared to Egoexo-Fitness Li et al. (2024) (see in Figure 9 and Table 2).
@@
Egoexo-Fitness Li et al. (2024) 1. Kneeling Push-ups 2. Push-ups 1. Pectoralis major 2. Anterior deltoid
@@ 13. Hip abductors
B.1.2 Subject Recruitment
Humans, as the core component in action performance, directly influence action quality. To collect more comprehensive data, the FLEX dataset required subjects across various capability levels compared with datasets that only contained expert-level subjects. So, we extensively recruited subjects within our institution and local commercial gyms, and ultimately selected 38 subjects, comprising 10 professional coaches, 8 amateurs, and 20 novices. We entered into a data-collection agreement with each subject, securing their authorization to record video and physiological data and to make these publicly available for academic research; the agreement also explicitly delineates the applicable financial compensation terms.
B.2 Collection
Instruction. As noted above, prior to data collection, the experimenter outfitted each subject with a custom vest and had surface EMG sensors affixed to the target muscle regions. Subjects then entered the capture area individually, and upon the experimenter’s start signal, each subject performed the action ten times based on their own experience. Notably, the experimenter informed subjects of the specific action to be recorded before each trial, and the equipment load was set to 80% of the subject’s maximum to ensure that the action was executed by their own capability.
Human subject safety. To ensure the safety of human subjects during data collection of risk-prone exercises, they were asked to perform exercises at 80% of their maximum weight-loaded capacity. Additionally, all subjects were covered by sports injury insurance, with on-site safety supervisors and emergency medical kits available throughout the process. Our data collection process was reviewed and approved by the Institutional Review Board (IRB). All participants provided written informed consent prior to participation.
Rest periods in between data collection. Considering the difficulty of performing 20 consecutive actions, particularly for the 20 novice subjects, only expert-level subjects performed all 20 consecutively. The remaining 28 subjects completed their collections separately over multiple days to ensure high-quality data, grouped by targeted muscle regions.
Given the differences in testing intensity, we implemented a differentiated compensation scheme: expert subjects received compensation equivalent to approximately 15 times the local minimum hourly wage for completing all tests within approximately two hours, while the other subjects completed their sessions across three days, about 30 minutes per day, and were compensated at a rate equivalent to roughly 7 times the local minimum hourly wage per day.
Motion capture environment and sensor placements. All devices were installed within a controlled space measuring $\times5m$$\times$2m (as shown in Figure 3). The four cinema cameras for motion capture were fixed at each corner of the square space, while the smartphone was positioned at the main viewpoint. All four sides of the controlled space were covered with curtains to reduce environmental interference further. The placement of the sEMG sensors was adjusted based on the specific target muscles involved in each action. Each subject wore the customized vest tightly to ensure accuracy in data acquisition.
B.3 Annotation
B.3.1 Annotation Rules
@@
Figure 10: The overview of the FLEX knowledge graph. (a) Visualization of frequently used annotation words. (b) FLEX-KG: the structure of the knowledge graph. (c1) Mapping between actions and action keysteps. (c2) Mapping between action keysteps and error types.
From the above (Sec. 3.3), we can see that the dataset annotation standards are largely derived from Chinese sources; however, they are by no means limited to one country. These standards are deeply informed by internationally recognized training principles and sports science knowledge. For instance, the curricula of Beijing Sport University and the guidelines from the General Administration of Sport of China are regularly benchmarked against frameworks from the American College of Sports Medicine (ACSM), the National Strength and Conditioning Association (NSCA), German sport science and rehabilitation models, Russian strength and conditioning traditions, and Australian applied sports science practices. Such integration ensures that the standards encompass comprehensive guidance on movement mechanics, posture, and safety considerations, making them broadly representative and internationally applicable rather than narrowly country-specific.
B.3.2 Annotator Recruitment
Due to the vast volume of the FLEX dataset and stringent annotation quality requirements, we recruited 16 professional practitioners from the fitness domain, including fitness trainers and graduate students in sports science, to participate in the annotation work. To ensure consistency and scientific rigor throughout the annotation process, we engaged in in-depth collaboration with all annotators before project commencement, clarifying the annotation requirements, personnel qualifications, and the underlying logic of the guidelines. To validate the feasibility and authority of these rules, we initially conducted a small-scale pilot annotation, during which we reviewed the annotators’ professional credentials and, based on the pilot results, made necessary revisions to the guidelines and refined the selection of annotators. The payment of a single annotator is higher than the local minimum hourly wage.
B.3.3 Annotation Consistency
As detailed in Sec. 3.3 and the Appendix B, we recruited 16 annotators and divided them into two groups according to experience level, employing a two-stage annotation process to ensure data quality. Specifically, in Stage 1, each sample was independently annotated for error types by three annotators from the first group, using both video and 3D pose data. This resulted in three initial annotations per sample. In Stage 2, a second group of three different annotators reviewed these annotations. For each sample, they assessed the initial labels and provided their own judgments. If disagreements were identified (i.e., inconsistent annotations or ambiguities), the second group performed a majority vote among themselves to resolve the conflict and produce a single final annotation. This process was designed to mitigate subjectivity and label noise. Ambiguous cases are discussed until an agreement is reached. For rare cases, authors invited domain experts from the Sports University to join meetings and facilitate consensus.
B.3.4 Score Calculation
During the annotation‐rule formulation phase, we assigned different penalty weights to error types according to their impact on action quality. We then summed the weights of all error types present in a given action to obtain that action’s total weight. For each sample, based on the error types identified by the annotators, we calculated the ratio of the sample’s cumulative error weight to the action’s total weight, and subsequently computed the action score using the following formula (see in Figure 11):
@@ B.3.4 Score Calculation
B.3.5 FLEX-VideoQA Dataset
We designed a dialogue template following the pipeline “action recognition → action standards → action evaluation → action scoring,” with all questions and reference answers automatically generated from our annotation rules and results. In particular, action-evaluation answers were pre-generated by DeepseekV3 by combining video samples, action keysteps, error types, and feedback suggestions to ensure consistency. Based on this pipeline, we constructed the FLEX-VideoQA dataset, which provides large-scale, fine-grained question–answer pairs tailored to action quality assessment.
Compared with existing non-AQA VideoQA tasks, FLEX-VideoQA is positioned with a clear distinction. Traditional non-AQA VideoQA datasets primarily focus on scene understanding, environmental context, or coarse-grained human actions, such as asking what a person is doing at a certain moment in the video or predicting what they might do next. Some datasets extend slightly further to include questions about more specific aspects, like the direction of a person’s gaze. However, these tasks remain largely at a high-level and coarse granularity. In contrast, FLEX-VideoQA targets the fine-grained aspects of actions and exercises, with particular emphasis on identifying errors in posture and movement. More importantly, it incorporates queries that require corrective feedback—for instance, what adjustments or improvements a person should make. Such queries, which shift the focus from simply recognizing “what is happening” to providing guidance on “how it should be improved,” are not addressed in existing non-AQA VideoQA datasets. This highlights the uniqueness and novelty of FLEX-VideoQA.
Figure 12: The construction of FLEX-VideoQA dataset.
B.4 Data Statistics